From nch@roe.ac.uk Fri Dec 14 11:33:11 2007 Date: Fri, 14 Dec 2007 11:27:24 +0000 (GMT) From: Nigel Hambly To: WFCAM Science Archive Team -- Eckhard Sutorius , Johann Bryant , Mike Read , Nigel Hambly , Nicholas Cross , Bob Mann , Ross Collins Cc: CCs for WSA weekly meeting minutes distribution -- Andrew Lawrence , Andy Adamson , Brian Walshe , Jim Emerson , Malcolm Stewart , Lorenzo Rimoldini , Mike Irwin , Mark Holliman , Peredur Williams , Stephen Warren Subject: WFAU VDFS Science Archive weekly project meeting minutes, 14/12/07 Minutes of WFAU VDFS Science Archive meeting: 14th December 2007 ------------------------------------------------------------------------- ------------------------------------------------------------------------- Present: NCH, RSC, MAR, JB, ETWS, LGR, NJC Apologies: BCW, JPE, AL, RGM, JMS, MSH, PMW DONM: 10am, Friday 11th January 2008 NB: in the VISTA HUT Actions discharged this week: ----------------------------- ACTION: JB to complete operations "brain dump" onto the TWiki. Discharged; today is JB's last day. ACTION: ETWS to put FreeTDS DNS entries in to enable ODBC/JDBC connectivity to hatshepsut. Discharged ACTION: NCH, ETWS, LGR and NJC to review their UKIDSS Workshop contributions next Wednesday at 2pm. Discharged Actions partly discharged but continuing: ----------------------------------------- The following from last time partly done but continue: o) Add in a default row for every detector appearing in every detection table (for schema consistency when querying merged sources and individual detections) - ACTION: RSC & NCH Continues; will get sorted (honestly!) as part of the general ingest DB schema revamp to facilitate rapid generalised photometric/astrometric recalibration after DR3. ACTION: MSH to bring hatshepsut onto private network, and to set up a login for MAR hatshepsut now on the archive private network; logins (both user and SQL) need to be fettled. Actions carried forward from 07/12/07 meeting: ---------------------------------------------- ACTION: JB & ETWS to restore SegueDR6 on thutmose at their convenience (i.e. no hurry given the present circumstances) Continues; team agreed to postpone until after DR3 now; also awaits instance of SQL 2005 on the batch catalogue server which in turn awaits full SSA tests against user interface JDBC applications... we're getting there... ACTION: JB to set dfg_1 shrinking in the WSA this weekend. Continues; some uncertainty as to whether this can be done during GPS reseaming; NCH to investigate. Specific points and new actions: -------------------------------- Project management: NCH apologised for any confusion over future involvement of JMS in VDFS activities; JMS noted: "The word "resignation" perhaps implies something sudden and unexepected... I won't be disappearing from VISTA though, since I'll continue to provide my services wearing a different hat, so not too much should change in the short term." WFCAM & VISTA updates: NCH noted that the JAC/CASU/WFAU MoU has been circulated amongst the UKIDSS survey heads; the team noted also the current financial climate and potential impact on UKIRT operations. Comments and issues arising from CASU fortnightly minutes: The team noted the minutes of the meeting of 11th December; there were no substantive comments. Networking: ETWS noted that the weird FreeTDS bug had reared it's ugly head again this week, so khafre's tmp partition filled up with SQL logs and froze that machine during GPS reseaming. JB noted some of the bulk data juggling that has been necessary this week; NCH asked about the old SSS dataset transfers and JB noted these had stalled over ftp installation on a NAS box. ACTION: NCH to buzz MSH about getting the SSS transfers instigated before Xmas. WSA Operations: GPS source merging finished earlier in the week; RSC noted that it took on average 1day to merge each semester - quicker than for DR2. Seaming is currently running and may finish before Xmas... ETWS, NCH and RSC to keep an eye on things over the next week. Hardware NCH thanked JB and MSH (in absentia) for their sterling efforts in rearranging all the archive hardware last weekend during the power outage. Software: RSC reported tracking down some obscure bugs this week as well as fixing some weird behaviour by CU7. Other than that, everyone has mucked in on the Plate Library removals. Survey Data Release: GPS for DR3 is continuing... Non-survey Data Release: Nothing to report this week Astrogrid deployment: NCH noted that *another* VO data centre questionnaire is being circulated: http://cds.u-strasbg.fr/twikiDCA/bin/view/EuroVODCA/DCACensusQuestionnaires but we await advice from the VO team here as to whether further inputs from the WFAU data holdings are necessary. Miscellaneous: Late-taste final warning for Fri 14th, Cannon's Gait from 8pm(ish) onwards for an Xmas p***-up featuring The Blackford Jazz Experience (good grief...)